# Adapting Noise to Data: Generative Flows from 1D Processes

### v1

@@ 1 Introduction
Flow-based generative models, especially score-based diffusion [30, 32], flow matching (FM) [1, 20, 21] and consistency models like the recently introduced inductive moment matching (IMM) [39], achieve state-of-the-art results in many applications. All these methods construct a probability flow from a simple latent distribution (noise) to a complex target (data) with a neural network trained to approximate this flow from limited target samples. In diffusion models, the score function directs a reverse-time SDE, while in FM, the velocity field is learned to compute trajectories via a flow ODE. Finally, consistency models like IMM learn to predict the jumps from noise to the data while factoring in the consistency of the flow trajectories. Usually, a Gaussian is used as latent distribution which causes difficulties when learning certain multimodal and heavy-tailed targets [13, 27], see Figure 5 for a heavy-tailed example. There exist only few approaches to learn the noising process, [3] fit the forward diffusion process via a learned invertible map that is trained end-to-end, [18] use metric flow matching, i.e., a neural network to adapt the path to a underlying Riemannian metric. In a related approach, [26] learns a componentwise Gaussian noise schedule, input-conditioned so each component receives its own noise level. On the other hand [23, 38] design heavy-tailed diffusions using Student-$t$ latent distributions, and [29] extend the framework to the family of $\alpha$-stable distributions.
@@ 4.2 Learning Quantile Processes
$\mathcal{L}(\theta,\phi)=\mathcal{L}_{\mathsf{OT-CFM}}(\theta,\phi)+\lambda\,\mathcal{L}_{\mathrm{AN}}(\phi),\qquad\lambda>0,$ (41)
@@
$\mathcal{L}_{\mathsf{OT-CFM}}(\theta,\phi)=\mathbb{E}_{t\sim\mathcal{U}(0,1),\ (x,y)\sim\pi_{\phi}}\Big[\big\|v_{\theta}\big((1-t)x+ty,\,t\big)-(y-x)\big\|_{2}^{2}\Big],$
@@
Algorithm 1 Joint learning of 1D quantiles and FM velocity
1:Dataset $\mathcal{D}$, batch size $B$, weight $\lambda_{Q}$, iterations $K$
@@
4:  Sample $\{\mathbf{x}_{i}\}_{i=1}^{B}\sim\mathcal{D}$, $\{\mathbf{u}_{i}\}_{i=1}^{B}\sim\mathcal{U}([0,1]^{d})$, $\{t_{i}\}_{i=1}^{B}\sim\mathcal{U}(0,1)$
5:  $C_{ij}\leftarrow\|\mathbf{x}_{i}-\mathbf{Q}_{\phi}(\mathbf{u}_{j})\|_{2}^{2}$
@@
7:  ${\mathcal{L}}_{\mathrm{AN}}\leftarrow\frac{1}{B}\sum_{i=1}^{B}\|\mathbf{x}_{i}-{\mathbf{Q}_{\phi}(u_{T(i)})}\|_{2}^{2}$
8:  $\mathbf{z}_{i}\leftarrow(1-t_{i})\mathbf{x}_{i}+t_{i}\,{\mathbf{Q}_{\phi}(u_{T(i)})}$
9:  ${\mathcal{L}_{\mathrm{OT-CFM}}}\leftarrow\frac{1}{B}\sum_{i=1}^{B}\|v_{\theta}(\mathbf{z}_{i},t_{i})-{\mathbf{Q}_{\phi}(u_{T(i)})}+\mathbf{x}_{i}\|_{2}^{2}$
10:  $\mathcal{L}\leftarrow{\mathcal{L}_{\mathrm{OT-CFM}}+\lambda\mathcal{L}_{\mathrm{AN}}}$
11:  Update $(\theta,\phi)$ by a gradient step on $L$
12:end for
13:return $(\theta,\phi)$
@@ 5.2 Image Datasets
Next we analyze our method on standard image generation benchmarks. Our quantile is extremely lightweight compared to the UNet architecture used for the flow model. We reuse the minibatch OT coupling for the latent and freeze the quantile function after a $25k$ training epochs. This strategy introduces only minimal computational overhead compared to the standard Gaussian baseline with minibatch OT coupling. Specifically, we observe an overhead of approximately $20\%$ in runtime during joint quantile training, and about $12\%$ after freezing the quantile parameters, measured on an NVIDIA GeForce RTX 4090.
@@ Appendix A Uniform Process and MMD Gradient Flow
fulfills $\mathcal{F}_{\nu}(\mu)=F_{\nu}(Q_{\mu})$ for all $\mu\in\mathcal{P}_{2}(\mathbb{R})$. Moreover, we have the following equivalent characterization of Wasserstein gradient flows of $\mathcal{F}_{\nu}$, which can be found in [9, Theorem 4.5].
@@ Appendix C IMM with Quantile Interpolants
Then it is proposed in [39, Eq. (7)] to minimize the so-called naïve objective
@@
Thus, for the number of steps $n\to\infty$, it holds $0={\rm MMD}^{2}_{K}(\rho_{s},p^{\theta_{n}}(s,t_{n}))$ even for the entire interval $t_{n}\in[s,1]$. Hence, minimizing the general objective (79) with a large number of steps eventually minimizes the naïve objective (78), see [39, Theorem 1] for more details.
@@ Appendix E Implementation Details
The coupling plans are calculated using the Python Optimal Transport package [11]. For inference simulate the corresponding ODEs using the torchdiffeq [5] package. For all models we only used the batch size $128$ and learning rate 2e42e-4 for the velocities. Quantile transports are parameterised by stacked rational-quadratic splines following [10].
@@ E.1 Synthetic Examples
For all models we used 3 hidden layers with width 64. We used a batch size of 128, a learning rate of 2e42e-4 and exponential moving average on the network weights of $0.999$. The baselines were trained for 200,000 iterations. We pretrain our quantiles and use the frozen quantiles during flow matching. We trained our quantile for 50,000 steps and to compensate we trained our velocity for only 150,000 steps. For the RQS we choose logit activation, 32 bins, a bound of 500 and one layer.
@@ E.2 Image Experiments
For the MNIST dataset we use the U-Net with base width 64, channel multipliers $(1,2,4)$, two residual blocks per resolution, attention at $7\times 7$, 1 attention head, and dropout 0.1. We clip the gradient norm to 1 and use exponential moving averaging with a decay of $0.99$. The quantiles were trained for the first 20,000 steps, after which the learning rate was linearly decayed to 0 by step 25,000.

### v2

@@ 1 Introduction
Flow-based generative models, especially score-based diffusion [32, 34], flow matching (FM) [1, 22, 23] and consistency models like the recently introduced inductive moment matching (IMM) [41], achieve state-of-the-art results in many applications. All these methods construct a probability flow from a simple latent distribution (noise) to a complex target (data) with a neural network trained to approximate this flow from limited target samples. In diffusion models, the score function directs a reverse-time SDE, while in FM, the velocity field is learned to compute trajectories via a flow ODE. Finally, consistency models like IMM learn to predict the jumps from noise to the data while factoring in the consistency of the flow trajectories. Usually, a Gaussian is used as latent distribution which causes difficulties when learning certain multimodal and heavy-tailed targets [15, 29], see Figure 5 for a heavy-tailed example. There exist only few approaches to learn the noising process, [3] fit the forward diffusion process via a learned invertible map that is trained end-to-end, [20] use metric flow matching, i.e., a neural network to adapt the path to a underlying Riemannian metric. In a related approach, [28] learns a componentwise Gaussian noise schedule, input-conditioned so each component receives its own noise level. In the setting of sampling from unnormalized target densities, [4] learn the latent noise by optimizing the mean and covariance of a Gaussian prior, while [5] learn a Gaussian mixture prior, both are trained end to end. On the other hand [25, 40] design heavy-tailed diffusions using Student-$t$ latent distributions, and [31] extend the framework to the family of $\alpha$-stable distributions.
@@ 4.2 Learning Quantile Processes
$\mathcal{L}(\theta,\phi)=\mathcal{L}_{\mathsf{OT\text{-}CFM}}(\theta,\phi)+\lambda\,\mathcal{L}_{\mathrm{AN}}(\phi),\qquad\lambda>0,$ (41)
@@
$\mathcal{L}_{\mathsf{OT\text{-}CFM}}(\theta,\phi)=\mathbb{E}_{t\sim\mathcal{U}(0,1),\ (x,y)\sim\pi_{\phi}}\Big[\big\|v_{\theta}\big((1-t)x+ty,\,t\big)-(y-x)\big\|_{2}^{2}\Big],$
@@
Algorithm 1 Joint learning of 1D quantiles and FM velocity (stop-gradient)
1:Dataset $\mathcal{D}$, batch size $B$, weight $\lambda$, iterations $K$
@@
4:  Sample $\{\mathbf{x}_{i}\}_{i=1}^{B}\sim\mathcal{D}$, $\{\mathbf{u}_{j}\}_{j=1}^{B}\sim\mathcal{U}([0,1]^{d})$, $\{t_{j}\}_{j=1}^{B}\sim\mathcal{U}(0,1)$
5:  $C_{ij}\leftarrow\|\mathbf{x}_{i}-\mathbf{Q}(\mathbf{u}_{j})\|_{2}^{2}$
@@
7:  Define $P$ by $P(j)=i$ such that $T(i)=j$
8:  $\hat{\mathbf{x}}_{j}\leftarrow\mathbf{x}_{P(j)}$
9:  $\mathbf{z}_{j}\leftarrow(1-t_{j})\hat{\mathbf{x}}_{j}+t_{j}\,\mathbf{Q}_{\phi}(\mathbf{u}_{j})$
10:  ${\mathcal{L}}_{\mathrm{AN}}\leftarrow\frac{1}{B}\sum_{j=1}^{B}\|\hat{\mathbf{x}}_{j}-\mathbf{Q}_{\phi}(\mathbf{u}_{j})\|_{2}^{2}$
11:  ${\mathcal{L}_{\mathrm{OT\text{-}CFM}}}\leftarrow\frac{1}{B}\sum_{j=1}^{B}\Big(\|v_{\theta}(\mathbf{z}_{j},t_{j})\|_{2}^{2}-2\langle v_{\theta}(\mathbf{z}_{j},t_{j}),\,\mathbf{Q}_{\phi}(\mathbf{u}_{j})-\hat{\mathbf{x}}_{j}\rangle+\|\mathbf{Q}(\mathbf{u}_{j})-\hat{\mathbf{x}}_{j}\|_{2}^{2}\Big)$
12:  $\mathcal{L}\leftarrow{\mathcal{L}_{\mathrm{OT\text{-}CFM}}+\lambda\mathcal{L}_{\mathrm{AN}}}$
13:  Update $(\theta,\phi)$ by a gradient step on $\mathcal{L}$
14:end for
15:return $(\theta,\phi)$
Remark 8.
In our implementation we optionally stop the gradient (w.r.t. $\phi$) through the pure $\|\mathbf{Q}_{\phi}(u)-x\|_{2}^{2}$ contribution inside the MSE $\|v_{\theta}(z,t)-(\mathbf{Q}_{\phi}(u)-x)\|_{2}^{2}$, while keeping gradients through the cross term and through $z=(1-t)x+t\mathbf{Q}_{\phi}(u)$. Concretely, we evaluate
$\|v_{\theta}(z,t)\|_{2}^{2}\;-\;2\langle v_{\theta}(z,t),\,\mathbf{Q}_{\phi}(u)-x\rangle\;+\;\|\mathbf{Q}(u)-x\|^{2},$
where the missing $\phi$ denotes stop–gradient. This detachment can slightly stabilize training, it is not necessary.
@@ 5.2 Image Datasets
Next we analyze our method on standard image generation benchmarks. Our quantile is extremely lightweight compared to the UNet architecture used for the flow model. We reuse the minibatch OT coupling for the latent and freeze the quantile function after a $25k$ training epochs. This strategy introduces only minimal computational overhead compared to the standard Gaussian baseline with minibatch OT coupling. On the CIFAR dataset for example, we observe an overhead of approximately $7\%$ in runtime during joint quantile training, and about $2\%$ after freezing the quantile parameters, measured on an NVIDIA GeForce RTX 4090.
@@ Appendix A Uniform Process and MMD Gradient Flow
fulfills $\mathcal{F}_{\nu}(\mu)=F_{\nu}(Q_{\mu})$ for all $\mu\in\mathcal{P}_{2}(\mathbb{R})$. Moreover, we have the following equivalent characterization of Wasserstein gradient flows of $\mathcal{F}_{\nu}$, which can be found in [11, Theorem 4.5].
@@ Appendix C IMM with Quantile Interpolants
Then it is proposed in [41, Eq. (7)] to minimize the so-called naïve objective
@@
Thus, for the number of steps $n\to\infty$, it holds $0={\rm MMD}^{2}_{K}(\rho_{s},p^{\theta_{n}}(s,t_{n}))$ even for the entire interval $t_{n}\in[s,1]$. Hence, minimizing the general objective (79) with a large number of steps eventually minimizes the naïve objective (78), see [41, Theorem 1] for more details.
@@ Appendix E Implementation Details
The coupling plans are calculated using the Python Optimal Transport package [13]. For inference simulate the corresponding ODEs using the torchdiffeq [7] package. For all models we only used the batch size $128$ and learning rate 2e42e-4 for the velocities. Quantile transports are parameterised by stacked rational-quadratic splines as described in D.2, we set the minimum bin width and height to 1e31e-3 and the minimum slope to 1e51e-5.
@@ E.1 Synthetic Examples
For all models we used 3 hidden layers with width 64. We used a batch size of 128, a learning rate of 2e42e-4 and exponential moving average on the network weights of $0.999$. The baselines were trained for 200,000 iterations. We pretrain our quantiles and use the frozen quantiles during flow matching. We trained our quantile for 50,000 steps and to compensate we trained our velocity for only 100,000 steps. Note however we still train our method For the RQS we choose logit activation, 32 bins, a bound of 500 and one layer.
@@ E.2 Image Experiments
For the MNIST dataset we use the U-Net with base width 64, channel multipliers $(1,2,4)$, two residual blocks per resolution, attention at $7\times 7$, 1 attention head, and dropout 0.1. We clip the gradient norm to 1 and use exponential moving averaging with a decay of $0.99$. The quantiles were trained for the first 20,000 steps, after which the learning rate was linearly decayed to 0 by step 30,000.